To neutralize 100 g of HNO3 solution, 30 cm3 of 15% NaOH solution with a density of 1.15 g / cm3 was used. Calculate the mass fraction (in%) of nitric acid in the solution!


1
Expert's answer
2021-05-13T02:09:25-0400

15% of NaOH means

In 100g solution 15g of NaOH

d = m/v

v = m/d = 100/1.15 = 86.95mL

Molarity of NaOH = "\frac{15}{40} × \frac{1000}{86.95}"

=4.31 M

moles of NaOH required = 4.31× "\frac{30}{1000}"

= 0.129moles

moles of HNO3 needed = 0.129 moles

"\frac{mass}{molar mass}" = 0.129

mass = 0.129 × molar mass

mass = 0.129×63 = 8.127g

mass fraction = "\frac{8.127}{100} × 100" = 8.127 %